A wire that is 76 inches lonh is bent to form a rectangle. If the length of this rectangle is 7 inches less than twice its width, then what is the area of the rectangle?

Answer by booklover15(10)

Length = Width - 7
L = W-7
76 = 2 ( L+W)
76/2 = L + W
38 = L + W
38 = W - 7 + W
38 + 7 = 2W
45 = 2W
W = 45/2
W = 22.5
L = W -7
L = 22.5 - 7 = 15.5
A = 15.5 x 22.5 = 348. 75